A Complete Bibliography of Publications in Applied Mathematics and Computation: 2025–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 December 2024 Version 1.07

Title word cross-reference

, 4 [LX25]. 1 [JSZ25]. 1n [AA25]. 2 [JCHW25, WHCB25]. 3 [MR25, ZHF25]. 4 [LZ25]. [LX25]. α [GCV25]. B [CCLL25]. D^{L} [ZWF25]. D^{Q} [ZWF25]. D_{2n} × Z_{m} [WZY25]. F [LX25]. H_{∞} [WZJW25]. L^{p} [CDM25]. M [AA25]. Q [ZZ25a]. N [SGH^{+}25, WWLZ25, ZWF25]. r [YZM25]. S [BFH^{+}25]. σ [KSFD25].

-Cayley [ZWF25]. -Edge [WHCB25]. -equivalent [CCLL25]. -factorization [JSZ25]. -fractal [GCV25]. -independence [LZ25]. -integral [ZWF25]. -irregularity [KSFD25]. -packing [BFH^{+}25]. -player [WWLZ25, SGH^{+}25]. -restricted [YZM25]. -saturating [LX25]. -spectral [ZZ25a]. -structure [YZM25]. -type [CDM25]. -walk-regular [ZHF25]. -Wright [AA25].

0/1 [ZZL^{+}25].

1 [Ano25a, Ano25b, Ano25c, Ano25d]. 15 [Ano25e, Ano25f, Ano25g, Ano25h].

2025 [Ano25a, Ano25e, Ano25b, Ano25f, Ano25c, Ano25g, Ano25h, Ano25d].

395 [AS25].

accounting [LL25]. accuracy [RPD25]. Acoustics [ALC^{+}25]. action [MMNP25]. active [CGH^{+}25]. actuator [BAT25, HWC25, SNP25, WZJW25, ZZ25c]. Adaptive [ZZ25d, BAT25, LHZL25, LWS^{+}25, LYD25, WZJW25, ZLA25, ZLCL25]. ADI [JCHW25]. adjustment [TLSVGR25]. advection [PP25, dOKA25]. advection-diffusion [PP25]. affine [IJM25]. against [SNP25]. agent [LHZL25, YGY^{+}25, ZZ25c]. agents [Zha25]. algorithm [GX25, MCHB25, Pal25, YWJ^{+}25]. algorithms [KXL^{+}25, LTH^{+}25]. Allee [CMG25]. Allen [OGPFVV25, WWQX25]. allocation [RS25]. almost [CMG25]. alternative [IÖÜP25]. among [Zha25]. Analysis [FBK25, GCV25, TF25, ZLHS25, ALC^{+}25, ALV25, BSG^{+}25, BM25, CWWT25, KXL^{+}25, PTB25, RPS^{+}25, YY25, ZZT25]. Analytically [Ruj25]. anti [LMQZ25]. anti-Ramsey [LMQZ25]. antiregular [KSFD25]. aperiodic [WSX^{+}25]. Appl [AS25]. application [BM25, CjZjHbX25, LYLS25, WZJW25, ZJX25]. applications [Liu25, MWS25, WX25]. applied [TNL25]. approach [AAS25, DAG25, EST^{+}25, HZF^{+}25, LCLY25, LW25a, QX25b, WZP^{+}25, YWJ^{+}25]. approximate [HYK^{+}25]. approximately [YGY^{+}25]. approximation [FLLR25, IJM25, RPD25]. approximations [BM25]. April [Ano25a, Ano25e]. arbitrary [CWWT25]. artificial [AS21, AS25]. assessment [YZM25]. asset [QZ25]. assets [XZN^{+}25]. assimilation [JH25]. associative [SNP25]. asymmetric [GLHS25]. asymmetrically [PPN^{+}25]. asymptotic [WDMM25]. asynchronous [MCH25b]. Attack [MCH25b]. Attack-compensated [MCH25b]. attacks [BAT25, WZJW25, YZ25, ZXW^{+}25, ZLCL25]. attenuated [MH25]. attitude [ZLCL25]. automata [WMW^{+}25]. avoidance [ZLA25].

B [PP25]. B-splines [PP25]. background [RS25]. Barrier [ZLA25, YWJ^{+}25]. Barrier-function [ZLA25]. based [CG25, CWWT25, DWL^{+}25, EST^{+}25, HYK^{+}25, HWC25, HCCH25, IÖÜP25, JYLS25, LHZL25, LWS^{+}25, LM25, PP25, Pal25, QX25a, RB25, TF25, WSX^{+}25, WWQX25, WXD^{+}25, YBSC25, YZM25, YZ25, ZWQC25, ZLA25, ZWL25, ZZL^{+}25]. basis [TNL25]. bead [MBF25]. bead-rod [MBF25]. behaviors [SGH^{+}25]. Bessel [ADZ25, KXL^{+}25]. between [QZ25]. bidirectional [SNP25]. bilinear [IJM25]. biological [LM25]. bipartite [SHY25, YT25]. blind [IAK25]. blood [RPS^{+}25]. Board [Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p]. Boltzmann [YQYS25]. Boolean [ALV25]. Boros [DS25]. boundary [ALC^{+}25, GCV25, HRS25, OGPFVV25]. Bounds [ZZ25a]. branch [BVZ25]. Bratu [SSH25]. Brownian [MBF25]. Bumpless [WZP^{+}25]. bundle [HZF^{+}25]. burning [LSQ25].

Cahn [OGPFVV25, WWQX25]. calculus [LYLS25]. capital [TLSVGR25]. capped [Ruj25]. capped/floored [Ruj25]. cardinality [BFH^{+}25]. case [MH25, Ruj25]. Cayley [WZY25, ZWF25]. cell [BM25, LM25]. cellular [WMW^{+}25]. center [LM25, WHCB25]. center-based [LM25]. chain [ZZT25]. chains [MBF25, ZW25]. Change [HCCH25]. Characterization [WHCB25]. Chebyshev [GKM25]. chimeras [PPN^{+}25]. class [BAT25, WDMM25, Wu25]. classes [ZW25]. closed [DZY^{+}25]. closed-loop [DZY^{+}25]. co [ZCJL25]. co-integrated [ZCJL25]. cobweb [LYLS25]. codes [Kov25]. coefficients [CWWT25, QXZ25, TO25, YQYS25]. coevolutionary [GPH25]. Cohen [ZN25]. collective [ZCJL25]. collision [ZLA25]. collocation [JLZ25, PP25, PTB25]. colorings [BFH^{+}25]. combination [Liu25]. combinatorial [ALV25]. Command [YBSC25]. Command-filter-based [YBSC25]. common [GPH25]. communication [WSX^{+}25, ZXW^{+}25, ZZ25c]. compact [MR25]. compensated [MCH25b]. compensation [GPH25]. competitive [FBK25]. complementarity [PSM25]. complements [AAV25]. complete [JSZ25, LYM25]. completion [HYK^{+}25]. Component [ZLW25]. components [BQS25]. Compt [AS25]. computation [BQS25]. concave [YWJ^{+}25]. conditions [ALC^{+}25, GCV25, KL25, Liu25, MM25, OGPFVV25]. configuration [MBF25]. conformal [HRS25]. conical [MH25]. conjecture [DS25]. conjugate [CFS25]. connected [lCXZ^{+}25]. connectedness [WHCB25]. connectivity [ZLW25]. consensus [GPL25, ZXW^{+}25, ZZ25c]. conservation [CG25, KBM25, TNL25]. considering [WSX^{+}25]. constrained [GPL25]. constraint [HWC25, Wu25]. constraints [LWS^{+}25]. contact [CjZjHbX25]. Contagion [KK25a]. contribution [ZCJL25]. control [BAT25, CGH^{+}25, GX25, GPL25, HWC25, IJM25, JYLS25, KNKR25, LHZL25, LWS^{+}25, LC25, MCH25b, MLTL25, QX25a, SNP25, WDMM25, WWS^{+}25, WZJW25, WZP^{+}25, YBSC25, YGY^{+}25, YZ25, ZWQC25, ZLA25, ZZ25c, ZLHS25, ZLCL25]. control-affine [IJM25]. convection [YQYS25]. convection-diffusion [YQYS25]. Convergence [PTB25, YY25, BSG^{+}25, BM25, PSM25]. convex [YY25]. convexities [ADPS25]. Cooperation [GSL25, GLHS25, LL25, LW25b, YYW^{+}25, Zha25, ZWL25]. cooperative [MLTL25]. correlation [QZ25]. Corrigendum [AS25]. cospectral [LSWZ25]. cost [YWJ^{+}25]. costly [AJK25]. Coupled [TGS25, PPN^{+}25]. cover [LCL25, ZHL25]. covers [QM25]. cross [FBK25]. cross-immunity [FBK25]. crystal [QHZ25]. cyber [DZY^{+}25]. cyber-physical [DZY^{+}25]. cycle [LCL25]. cyclic [ZW25]. cylinder [MH25].

D [JCHW25, MR25]. Data

[IJM25, HZF^{+}25, JH25, TF25, WHCB25, WDMM25, WSX^{+}25, YZ25]. Data-driven [IJM25]. Davey [HAD25]. DCMIP [WZJW25]. DDSIIM [CWWT25]. dead [LHZL25]. dead-zone [LHZL25]. deadline [QZ25]. deblurring [IAK25]. deception [WZJW25, ZLCL25]. decision [ZJX25]. decision-making [ZJX25]. decisions [AJK25]. Decoupled [QHZ25]. Deep [AAS25, ZZL^{+}25]. Deep-time [AAS25]. defined [ZCJL25]. degenerate [KK25b, WZJW25]. delay [DZY^{+}25, PTB25, SGH^{+}25, WSX^{+}25, YZ25, ZZ25b]. delayed [CCLL25, SZY25, SNP25, WDMM25, WSX^{+}25]. delays [ZN25]. demand [ZLHS25]. dependent [BQS25, CCLL25, YQYS25]. derivative [LT25]. derivatives [AS21, AS25]. descent [YWJ^{+}25]. described [XZN^{+}25]. description [MMNP25]. detection [HCCH25]. determinant [Ued25, ZJX25]. Determination [JH25]. deterministic [DAG25]. device [WZJW25]. diagonalization [LCLY25]. diameter [LSWZ25, WJBL25]. diffeomorphism [LBLZ25]. difference [OGPFVV25, SSH25]. differential [CjZjHbX25, GPL25, QXZ25, XZN^{+}25]. diffusion [BM25, PP25, QX25b, Ruj25, YQYS25, ZWL25, dOKA25]. dilemma [LW25b, YYW^{+}25]. dilute [MBF25]. dimensional [AAS25, LJ25, SSH25, WMW^{+}25, WWS^{+}25]. direction [YWJ^{+}25]. directions [QX25a]. Dirichlet [ALC^{+}25]. discontinuous [CWWT25]. Discrete [DZY^{+}25, IJM25, JCHW25, LYLS25, Ruj25]. Discrete-time [DZY^{+}25, IJM25]. disease [FBK25]. Disjoint [QM25, LCL25]. dissipative [WWS^{+}25]. Distance [ZW25, BFH^{+}25]. Distinguishing [AAV25]. Distributed [LYD25, LC25, ZXW^{+}25, ZZ25c, ZZ25d]. distribution [MP25, QX25b]. Disturbance [LWS^{+}25, ZZ25b]. disturbances [CGH^{+}25, YGY^{+}25]. diverse [SGH^{+}25]. domain [MCH25a]. domains [CWWT25, TKM25]. domatic [GVW25]. domination [GVW25]. double [RKS25]. driven [lCXZ^{+}25, GPL25, IJM25, LW25b, Ruj25]. dual [MWS25]. due [DS25]. Dynamic [GPL25, ZLCL25, HWC25, QZ25, WZP^{+}25]. dynamical [CGH^{+}25]. dynamics [ALV25, CMG25, Che25, IJM25, LDT^{+}25, SGH^{+}25, ZWL25].

ecological [GPH25]. ecosystem [ZWL25]. Edge [WHCB25]. edges [LX25]. Editorial [Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p]. effect [CMG25, GPH25, MLTL25]. effects [FBK25, LYD25, SGH^{+}25]. Efficient [HRS25, LT25, TKM25, AAS25]. eigenvalue [LYM25, MP25, XSB25, ZHF25]. element [HAD25, LT25, LJ25, Tan25, TKM25]. elliptic [CWWT25]. endogenous [GSL25]. energies [ZW25]. enhanced [LT25]. Enhancement [LDT^{+}25]. entropy [CLLC25]. Epidemic [FBK25]. epidemics [MMNP25]. equality [LL25]. equation [GKM25, JCHW25, LTH^{+}25, PP25, QX25b, SSH25, WWQX25, YQYS25]. equations [ALC^{+}25, FLLR25, HAD25, JLZ25, MR25, OGPFVV25, PTB25, QXZ25, TGS25, Tan25, TKM25, XZN^{+}25]. equienergetic [LSWZ25]. equivalent [CCLL25, Liu25]. Erdos [MMNP25]. error [CWWT25, CLLC25, HAD25]. errors [ZJX25]. estimates [TNL25]. Estimation [XZN^{+}25, ZWQC25, EST^{+}25, HAD25, LYD25, SZY25, WXD^{+}25, dOKA25]. Estimation-based [ZWQC25]. ETF [WZJW25]. Euler [TKM25]. Event [QX25a, YGY^{+}25, ZZ25c, DZY^{+}25, GPL25, HWC25, LHZL25, LC25, WZP^{+}25, YBSC25, YZ25, ZWQC25, ZLCL25]. event-driven [GPL25]. Event-triggered [QX25a, YGY^{+}25, ZZ25c, DZY^{+}25, HWC25, LHZL25, LC25, YZ25, ZWQC25, ZLCL25]. event-triggering [WZP^{+}25, YBSC25]. Evolution [GLHS25, Zha25, DWL^{+}25, MCH25a, TF25]. evolutionary [LL25, SGH^{+}25, ZLHS25]. expansion [BSG^{+}25]. explicit [dOKA25]. Exploring [MMNP25, YT25, ZWL25]. Exponential [CCLL25, MR25, ZN25]. expressions [WWLZ25]. extend [CLLC25]. extinction [CMG25]. extrema [LZ25]. Extremizing [KSFD25].

factorization [JSZ25]. failure [SNP25].

fairness [DWL^{+}25, Zha25].

fairness-seeking [Zha25]. false [YZ25]. Fast [KXL^{+}25, Pal25, Tan25]. Fault [KNKR25, YZM25, CGH^{+}25, EST^{+}25]. Fault-tolerant [KNKR25, CGH^{+}25]. faults [BAT25, CGH^{+}25, HWC25, WZJW25, ZZ25c]. fear [MLTL25]. February [Ano25b, Ano25f]. feedback [BAT25, GX25, MCH25b, SNP25]. FFT [CWWT25]. FFT-based [CWWT25]. fiducial [CLLC25]. field [MMNP25, QHZ25, ZZT25]. fields [MBF25]. Filippov [ZZT25]. filter [CLLC25, YBSC25]. filtering [BGQR25]. Finite [IJM25, WWS^{+}25, CF25, EST^{+}25, HWC25, LT25, LJ25, OGPFVV25, SSH25, Tan25, WMW^{+}25, ZJX25]. finite-difference [OGPFVV25]. Finite-region [WWS^{+}25]. finite-time [EST^{+}25, HWC25]. first [GKM25]. Fixed [CGH^{+}25, QX25a, ZWQC25]. Fixed-time [CGH^{+}25, QX25a, ZWQC25]. flight [KNKR25]. floored [Ruj25]. flow [MBF25]. flows [lCXZ^{+}25]. fluid [KBM25]. fluid-poroelastic [KBM25]. food [ZZT25]. forces [LM25]. form [KL25, MR25, TGS25]. formation [KNKR25, YGY^{+}25]. forms [AAV25, LBLZ25]. formulae [WX25]. formulas [MH25]. four [LSWZ25]. Fourier [CDM25]. fourth [PP25]. fractal [GCV25]. fractional [BGI25, CjZjHbX25, CDM25, JCHW25, JLZ25, LYLS25, QX25b, SHY25, Tan25, WWQX25, WXD^{+}25, XZN^{+}25, ZN25]. fractional-order [WXD^{+}25]. fragile [SNP25]. framework [MAMFBM^{+}25]. Fredholm [QXZ25]. free [LT25, QX25a]. frictional [CjZjHbX25]. frictions [RS25]. full [LWS^{+}25]. Fully [LC25, ZXW^{+}25, JCHW25]. function [ADZ25, AA25, TNL25, ZLA25, ZZL^{+}25]. functions [GCV25, MCHB25]. Fuzzy [LYLS25, EST^{+}25, KL25, LWS^{+}25, ZWQC25, ZZ25b].

game [GSL25, GPH25, LW25b, SGH^{+}25, YYW^{+}25, Zha25, ZLHS25]. games [GPL25, IÖÜP25, LL25, Ued25, WWLZ25, ZJX25]. gasket [GCV25]. gauge [SHY25]. general [ADPS25, RKS25, TGS25, TO25, Wu25]. generalized [HZF^{+}25, KXL^{+}25, LZ25, LJ25, Wu25, YY25, ZHL25]. genuine [ZHF25]. geostrophic [BGQR25]. given [LZ25]. GL-ADI [JCHW25]. glioma [BM25]. Global [WDMM25, FLLR25]. glucose [AS21, AS25]. glucose-insulin [AS21]. Godunov [RPD25]. Godunov-type [RPD25]. goods [GSL25, Zha25]. gradient [CFS25]. granular [LYLS25]. graph [ADPS25, HCCH25]. graphical [GPL25]. Graphs [LSQ25, AAV25, BFH^{+}25, CDD^{+}25, GVW25, KSFD25, LYM25, LZ25, MP25, QM25, RKS25, WZY25, WMW^{+}25, WJBL25, YT25, YZM25, ZHF25, ZWF25]. grid [Tan25]. grids [ZLHS25]. Grossberg [ZN25]. group [ZJX25]. growth [BDL25]. guidance [YYW^{+}25].

H [DZY^{+}25]. H-infinity [DZY^{+}25]. Hamiltonian [WHCB25]. Hammerstein [WXD^{+}25]. Hamming [YZM25]. hand [CF25]. harmonic [KK25b]. Heisenberg [CDM25]. hemivariational [CjZjHbX25]. Hermite [ADZ25, WWQX25]. hesitancy [LdMAS25]. heterogeneous [ZWL25, ZZ25c]. hidden [LW25a]. high [AAS25, CLLC25, LT25, MR25, QXZ25, SSH25, ZWQC25, ZLA25]. high-dimensional [AAS25, SSH25]. high-order [CLLC25, LT25, MR25, QXZ25, ZWQC25, ZLA25]. higher [KK25b, LDT^{+}25]. higher-order [KK25b, LDT^{+}25]. Hilbert [Liu25, QXZ25]. Hilbert-type [Liu25]. holistic [IÖÜP25]. homogeneous [LJ25]. horizon [LYD25]. human [lCXZ^{+}25, LW25a, YYW^{+}25]. human-driven [lCXZ^{+}25]. human-machine [LW25a]. hunting [MLTL25, ZWL25]. Hybrid [MLTL25]. hydrodynamics [RPD25]. hyperbolic [CMG25, PP25]. hypercube [CDD^{+}25]. hypercube-like [CDD^{+}25]. hypercubes [KV25, LMQZ25]. hypergraph [JSZ25]. hypergraphs [MAMFBM^{+}25, ZZ25a]. hypersingular [GKM25].

identification [QX25b]. identities [KK25b]. image [IAK25]. immobile [Tan25]. immunity [FBK25]. Impenetrability [ALC^{+}25]. imperfect [LdMAS25]. implementation [Ued25, ZJX25]. implicit [MR25]. Improved [MCHB25]. Improving [RPD25, ZZ25b]. impulses [CCLL25]. impulsive [CjZjHbX25, MM25, WZJW25, ZN25]. incidence [ZZ25a]. inclusions [CjZjHbX25]. incorporating [QHZ25]. independence [LZ25]. index [WJBL25]. indices [WZY25, ZHL25]. inequalities [HZF^{+}25]. inequality [KL25, WSX^{+}25]. inexact [HZF^{+}25]. infinite [CF25]. infinitesimal [CF25]. infinity [DZY^{+}25]. Influence [LdMAS25, DAG25]. initial [QX25b]. injection [BAT25, LBLZ25, YZ25]. input [QX25a, ZZ25d]. Insider [QZ25]. insulin [AS21, AS25]. insurable [RS25]. Integral [dOKA25, FLLR25, GKM25, Liu25, PTB25, TGS25, WSX^{+}25, ZWF25]. integrated [ZCJL25, ZZT25]. integration [LM25]. integro [QXZ25]. integro-differential [QXZ25]. interaction [KBM25]. interactions [LDT^{+}25, QHZ25]. interconnected [JYLS25]. interface [CWWT25, LJ25]. intermittent [CGH^{+}25]. interval [EST^{+}25, ZZ25b]. inverse [XSB25]. inverses [MWS25]. Inversion [MH25]. investment [RS25, SGH^{+}25]. irregular [MCH25a]. irregularity [KSFD25]. iteration [ADPS25].

Jacobi [JLZ25, RB25, XSB25]. January [Ano25c, Ano25g]. Jarratt [BSG^{+}25]. Jarratt-type [BSG^{+}25]. job [AJK25]. jump [JYLS25, Ruj25, WZJW25]. jump-diffusion [Ruj25]. jumping [WWS^{+}25].

Kalman [CLLC25]. kernel [Liu25, QXZ25, RB25]. kernels [PTB25]. kind [FLLR25, GKM25]. kinetics [AS21, AS25]. Kirchhoff [WZY25]. Koopman [IJM25]. KVA [TLSVGR25].

labelling [BVZ25]. labor [ZCJL25]. lag [MLTL25]. Lagrange [YWJ^{+}25]. Laguerre [ADZ25]. Lanzhou [WJBL25]. Laplacian [LYM25, MP25, ZW25]. lattice [YQYS25]. law [LTH^{+}25, MMNP25]. laws [CG25, TNL25]. layer [BGQR25]. learning [JYLS25, LHZL25, PPN^{+}25, YGY^{+}25, YZ25]. learning-based [JYLS25, LHZL25]. least [LYM25, ZHF25]. Level [IAK25, MR25]. Lexicographic [AS21, AS25]. like [CDD^{+}25, MM25]. limits [MMNP25]. line [WXD^{+}25]. Linear [BGQR25, CF25, CMG25, JYLS25, KK25a, KL25, MR25, QHZ25, QXZ25, Tan25, TO25, TNL25, Wu25, ZZ25d, dOKA25]. linear-hyperbolic [CMG25]. link [ZXW^{+}25]. liquidity [QZ25]. Local [LTH^{+}25, Pal25, PSM25]. logarithmic [YWJ^{+}25]. long [MCH25a, QHZ25]. long-range [QHZ25]. long-term [MCH25a]. loop [DZY^{+}25]. Lotka [Che25]. Low [HYK^{+}25]. Low-rank [HYK^{+}25]. Lur’e [HWC25]. Lyapunov [MM25]. Lyapunov-like [MM25].

Ma [DS25]. machine [LW25a]. maintains [YYW^{+}25]. making [ZJX25]. management [ZZT25]. manifolds [CFS25]. mappings [HRS25]. March [Ano25d, Ano25h]. markets [ZCJL25]. Markov [BM25, IÖÜP25, JYLS25, LW25a, WWS^{+}25]. Mass [KBM25, MMNP25]. matching [LMQZ25, ZHL25]. Math [AS25]. Mathematical [TLSVGR25]. matrices [MWS25]. Matrix [WWLZ25, HYK^{+}25, IÖÜP25, KL25, WSX^{+}25]. matrix-separation-based [WSX^{+}25]. Maximal [WJBL25]. mean [MMNP25]. measurements [WXD^{+}25]. measures [HCCH25]. mechanical [LM25]. mechanism [Che25, DWL^{+}25, WWS^{+}25, YBSC25]. memetic [Pal25]. memory [HWC25, SNP25, ZWL25, ZN25]. memory-based [HWC25, ZWL25]. memristive [PPN^{+}25, SZY25, SNP25]. Mental [LL25]. Merton [Ruj25]. meshless [MCH25a]. metastatic [BDL25]. Method [TO25, CFS25, CCLL25, FLLR25, HAD25, IAK25, IÖÜP25, JLZ25, LT25, MCH25a, PP25, PSM25, QXZ25, RB25, SSH25, TKM25, WXD^{+}25, YY25, ZZL^{+}25, ZLHS25]. methods [BSG^{+}25, GKM25, RPD25, Tan25, TNL25, TLSVGR25]. microgrids [LC25]. migration [BM25]. mimetic [OGPFVV25]. MIMO [LBLZ25]. Minimizing [LYM25]. Minimum [CLLC25, YWJ^{+}25]. mining [TF25]. mixed [HRS25]. mobile [Tan25]. mobile/immobile [Tan25]. mode [KNKR25]. model [BGQR25, CMG25, Che25, KK25a, KNKR25, MLTL25, RPS^{+}25, Ruj25, TGS25, YQYS25, ZZT25]. modeling [LW25a]. Modelling [DAG25]. models [BDL25, LYLS25, LM25, TLSVGR25]. modification [XSB25]. modified [QHZ25]. modifying [KSFD25]. Moll [DS25]. mollifier [QX25b]. Moore [MWS25]. most [LSWZ25]. moving [LYD25]. Multi [ALV25, KNKR25, LHZL25, LYD25, YGY^{+}25, ZZ25c]. multi-agent [LHZL25, YGY^{+}25, ZZ25c]. multi-sensor [LYD25]. Multi-state [ALV25]. multi-UAV [KNKR25]. multiagent [GPL25, ZXW^{+}25]. Multiple [ZN25, XZN^{+}25]. Multiscale [MBF25]. multispace [Kov25]. multispaces [Kov25]. multistate [BQS25]. multivariate [MCHB25]. Müntz [WX25]. Mutual [CDD^{+}25, RKS25, KV25]. Mutual-visibility [RKS25, KV25]. Mycielskians [RKS25].

Navier [JH25, JLZ25]. nested [KL25]. network [PPN^{+}25]. networked [YZ25]. networks [AAS25, ALV25, HCCH25, LCL25, LYD25, MMNP25, SZY25, SHY25, SNP25, WHCB25, WSX^{+}25, YBSC25, ZLW25, ZN25, ZZL^{+}25]. neural [AAS25, SZY25, SNP25, WSX^{+}25, WWQX25, YBSC25, ZN25, ZZL^{+}25]. neuron [PPN^{+}25]. neutral [YYW^{+}25]. Newton [BGI25, LT25, YY25]. Non [LSWZ25, SNP25, JSZ25, LJ25, Tan25, dOKA25]. Non-cospectral [LSWZ25]. Non-fragile [SNP25]. non-homogeneous [LJ25]. non-linear [Tan25, dOKA25]. non-uniform [JSZ25, Tan25]. nonhomogeneous [Ruj25]. Nonlinear [CFS25, LBLZ25, ALC^{+}25, BAT25, BGQR25, EST^{+}25, FLLR25, GX25, IJM25, LWS^{+}25, LTH^{+}25, PSM25, QX25a, RS25, Ruj25, TGS25, WXD^{+}25, YBSC25, ZWQC25, ZLA25, ZXW^{+}25]. nonlinearity [LTH^{+}25]. nonlocal [CjZjHbX25]. nonperiodic [OGPFVV25]. nonsmooth [AS21, AS25]. norm [IÖÜP25, Liu25]. norm-based [IÖÜP25]. normal [AAV25, AS21, AS25, LBLZ25]. Note [LMQZ25, JSZ25]. novel [MCH25a]. number [ADPS25, LSQ25, LMQZ25, LX25, LZ25]. numbers [GVW25, KK25b]. Numerical [ALC^{+}25, BM25, BDL25, HAD25, LM25, WWQX25, TLSVGR25].

observable [LBLZ25]. observation [Ruj25]. observer [EST^{+}25, LWS^{+}25, ZZ25d]. observer-based [EST^{+}25, LWS^{+}25]. ocean [BGQR25]. on-line [WXD^{+}25]. one [WMW^{+}25, XSB25]. one-dimensional [WMW^{+}25]. operator [IJM25, Liu25]. operators [OGPFVV25]. Optimal [AJK25, RS25, ZCJL25, GPL25, MCHB25]. optimization [CFS25, GX25, LCLY25, YY25]. optimized [YGY^{+}25]. options [AJK25, Ruj25, XZN^{+}25]. order [AA25, CG25, CLLC25, KK25b, LT25, LZ25, LDT^{+}25, MR25, PP25, QXZ25, SZY25, Tan25, WXD^{+}25, ZWQC25, ZLA25]. ordering [Pal25]. orthogonal [LCLY25]. oscillatory [KXL^{+}25]. output [BAT25, GX25, LBLZ25, MCH25b, SNP25, YZ25]. output-feedback [BAT25, SNP25].

packing [BFH^{+}25]. pancreatic

[AS21, AS25]. pancyclicity [LCL25]. paper [LDT^{+}25]. parabolic [EST^{+}25, MR25]. parallel [Pal25]. Parameter [WXD^{+}25, JH25]. parameterized [KL25]. parameters [XZN^{+}25]. partially [JYLS25]. particle [RPD25]. passivity [SHY25]. Past [LW25b]. Past-performance-driven [LW25b]. path [QM25]. Pattern [Che25, YZM25]. Pauli [CDM25]. payoff [LL25]. payoffs [Ruj25]. PDE [EST^{+}25]. PDEs [AAS25]. peers [LdMAS25]. Penrose [MWS25]. Perfect [ZHL25]. performance [HWC25, LW25b, ZZ25b]. periodic [OGPFVV25, WMW^{+}25]. persistence [LDT^{+}25]. perspective [LL25]. perturbed [MM25]. pest [ZZT25]. phase [QHZ25]. phenomenon [AA25]. physical [DZY^{+}25]. piecewise [BM25]. Pinning [SHY25]. Plane [MH25]. plans [ZCJL25]. player [SGH^{+}25, WWLZ25]. point [ADZ25, GCV25, HCCH25, Wu25]. points [CLLC25]. Poisson [Ruj25]. polymeric [MBF25]. polynomial [PP25, RB25]. polynomials [DS25, RB25, WX25]. polyomino [ZW25]. pool [GPH25]. population [CMG25]. populations [LM25, YYW^{+}25]. poroelastic [KBM25]. portfolio [AJK25, RS25]. position [ADPS25, RKS25]. positivity [QHZ25]. positivity-preserving [QHZ25]. poverty [DAG25]. power [LTH^{+}25]. Practical [HWC25]. preconditioners [Wu25]. predator [MLTL25, TGS25]. predator-prey [TGS25]. predators [ZWL25]. predefined [YBSC25, ZLCL25]. predefined-time [ZLCL25]. predictive [KNKR25]. Predictor [YZ25, IJM25]. Predictor-based [YZ25]. prescribed [MM25, WJBL25]. prescribed-time [MM25]. presence [BAT25]. preserving [LTH^{+}25, QHZ25]. prevalence [DAG25]. prey [MLTL25, TGS25]. prey-predator [MLTL25]. pricing [Ruj25]. principles [CDM25]. prisoner [LW25b, YYW^{+}25]. Probabilistic [LW25a]. probability [KK25a, MCH25b]. problem [CjZjHbX25, MCH25a, Pal25, TLSVGR25, XSB25, YWJ^{+}25]. problems [BGI25, CWWT25, HRS25, IAK25, KV25, LJ25, YY25, dOKA25]. procedure [BGI25]. process [Ruj25]. processes [BM25]. programming [CF25]. projection [GKM25]. promote [LW25b]. promotes [DWL^{+}25]. promotion [LL25]. properties [SSH25]. prostitution [DAG25]. proximal [HZF^{+}25]. pseudo [XSB25]. pseudo-Jacobi [XSB25]. public [GSL25, Zha25]. punishment [GSL25].

Qi [DS25]. quadratic [JYLS25]. Qualitative [ZZT25]. quantization [LYD25, MCHB25]. quantized [HWC25]. quarantine [FBK25]. quasi [BGQR25, MR25]. quasi-geostrophic [BGQR25]. quasi-linear [MR25]. quaternion [HYK^{+}25]. QUAV [ZLCL25].

radial [TNL25]. Radio [BVZ25]. radius [ZZ25a]. Radon [MH25]. Ramsey [LMQZ25]. random [MCHB25, QZ25]. range [QHZ25]. rank [HYK^{+}25, XSB25]. rank-one [XSB25]. Reachable [SZY25]. reaction [JCHW25]. reaction-subdiffusion [JCHW25]. recurrence [WX25]. recursion [ZZL^{+}25]. Redefined [PP25]. redundant [YT25]. region [WWS^{+}25]. regular [ZHF25]. regularity [Tan25]. regularization [YY25]. regularized [YY25, YQYS25]. regularizer [HYK^{+}25]. Reinforcement [JYLS25, LHZL25, YGY^{+}25]. rejection [ZZ25b]. relaxation [CG25, YQYS25]. Relaxed [KL25]. reliability [BQS25]. Rényi [MMNP25]. repeated [Ued25]. replanning [KNKR25]. Reproducing [QXZ25]. reputation [DWL^{+}25]. reputation-based [DWL^{+}25]. Resilient [EST^{+}25, LC25]. Resolving [WMW^{+}25]. resource [GPH25]. respiratory [FBK25]. response [ZLHS25]. responses [AS21, AS25]. restricted [YZM25]. results [GKM25, WHCB25]. retirement [AJK25]. retraction [CFS25]. reversibility [WMW^{+}25]. reward [DWL^{+}25, IÖÜP25]. rewards [GLHS25]. Riemannian [CFS25]. right [CF25]. right-hand [CF25]. risk [RS25]. robust [RPS^{+}25]. rock [LDT^{+}25]. rod [MBF25]. Roesser [WWS^{+}25]. Roman [GVW25]. rotation [ZHL25]. row [Pal25].

S [ZZ25b]. saddle [Wu25]. sampled [WDMM25, WSX^{+}25]. sampled-data [WDMM25, WSX^{+}25]. sampling [QX25a]. sampling-based [QX25a]. saturating [LX25]. saturation [QX25a]. scalar [CG25, TNL25]. scale [MAMFBM^{+}25]. scaling [ZXW^{+}25]. scheme [JCHW25, MR25]. schemes [CG25, QHZ25]. Schrödinger [LTH^{+}25]. scissors [LDT^{+}25]. search [Pal25]. secant [PSM25]. Second [CG25, Tan25, FLLR25, LT25, SZY25]. second-derivative-free [LT25]. second-kind [FLLR25]. Second-order [CG25, Tan25, SZY25]. secondary [LC25]. Secure [BAT25, ZXW^{+}25]. seeking [Zha25]. self [ZXW^{+}25]. self-triggered [ZXW^{+}25]. semi [LW25a]. semi-Markov [LW25a]. sensor [BAT25, LYD25]. separable [WXD^{+}25]. separation [WSX^{+}25]. sequence [GX25]. sequential [ALV25, ZXW^{+}25]. set [SZY25]. sets [BFH^{+}25]. shape [RPS^{+}25]. shooting [BGI25]. shooting-Newton [BGI25]. short [ZN25]. side [CF25]. Sierpiński [GCV25]. Signed [GVW25, LYM25, SHY25]. signless [ZW25]. similarity [HCCH25]. simplified [YGY^{+}25]. simulation [HRS25, HAD25, WWQX25]. simulations [MBF25]. simultaneous [LCLY25, QX25b]. single [MCHB25]. single- [MCHB25]. singular [JYLS25, KXL^{+}25, PTB25]. singularity [QX25a]. singularity-free [QX25a]. sliding [KNKR25, ZZT25]. slip [JH25]. smart [ZLHS25]. Smith [AAV25]. smoothed [RPD25]. smoothing [PSM25]. snapshot [HCCH25]. snarks [ZHL25]. social [LdMAS25]. sojourn [MCH25b]. solution [BDL25, IÖÜP25]. solutions [MBF25, PTB25, TO25]. solver [CWWT25, WWQX25]. solvers [KBM25]. Solving [OGPFVV25, AAS25, BGI25, HZF^{+}25, MCH25a, PP25]. Some [KK25b]. source [QX25b, dOKA25]. space [JLZ25, QX25b, QXZ25]. space-fractional [JLZ25, QX25b]. sparse [HYK^{+}25, LCLY25, TO25]. spatial [GSL25, LW25b, MLTL25, Zha25]. spatially [YQYS25, ZWL25]. spectra [AAV25, ZW25]. spectral [GKM25, JLZ25, LZ25, TKM25, ZZ25a]. splines [PP25]. split [LCL25]. split-star [LCL25]. Stability [TNL25, CCLL25, LW25a, MM25, ZN25]. stabilization [DZY^{+}25, LW25a, WDMM25]. Stable [LJ25]. star [LCL25, QM25]. state [ALV25, CCLL25, EST^{+}25, GPL25, LWS^{+}25]. state-constrained [GPL25]. state-dependent [CCLL25]. static [GX25]. step [ZZL^{+}25]. Stewartson [HAD25]. stochastic [DAG25, MCH25b, QZ25, TGS25]. stock [ZCJL25]. Stokes [AA25, JLZ25]. strategies [Ued25, ZCJL25]. strategy [HWC25, LT25, LW25b, ZJX25, ZZT25]. strongly [ZHF25]. structure [KBM25, LTH^{+}25, YZM25]. structure-preserving [LTH^{+}25]. structured [PSM25]. Study [lCXZ^{+}25]. subdiffusion [JCHW25]. subject [LYD25, QX25a]. substructure [YZM25]. sum [GPL25]. summations [KL25]. Superconvergence [GKM25]. supply [YWJ^{+}25]. sure [CMG25]. SVD [HYK^{+}25]. swaps [Ruj25]. switched [CCLL25, LWS^{+}25, WZP^{+}25, YBSC25]. switching [AJK25, GX25, MCH25b, WZP^{+}25]. symmetric [WWLZ25]. symmetry [SSH25]. Synchronization [PPN^{+}25, HWC25, SHY25, WSX^{+}25]. synthesis [GX25]. system [BQS25, CG25, DZY^{+}25, FBK25, LW25a, QHZ25, ZZ25c]. systems [BAT25, CG25, CGH^{+}25, CCLL25, EST^{+}25, GX25, HWC25, JYLS25, LHZL25, LWS^{+}25, MM25, MCH25b, QX25a, TO25, WDMM25, WXD^{+}25, WWS^{+}25, WZJW25, WZP^{+}25, Wu25, YBSC25, YGY^{+}25, YZ25, ZWQC25, ZLA25, ZXW^{+}25, ZZ25b, ZZ25d].

T [ZZ25b]. Tabu [PPN^{+}25]. taxis [Che25]. Taylor [BSG^{+}25]. temporal [HCCH25]. tensor [LCLY25]. term [MCH25a]. terminal [BGI25]. terms [YY25]. their [AAV25, BSG^{+}25]. Theory [ADZ25]. three [DZY^{+}25, LSQ25, LJ25]. three-dimensional [LJ25]. threshold [KK25a, ZWL25]. thresholds [FBK25]. Time [BQS25, SGH^{+}25, AAS25, ADPS25, CGH^{+}25, DZY^{+}25, EST^{+}25, HWC25, IJM25, JCHW25, LHZL25, MM25, MLTL25, QX25a, Tan25, WDMM25, WWQX25, YBSC25, YQYS25, ZWQC25, ZN25, ZZ25b, ZLCL25]. time-delayed [WDMM25]. Time-dependent [BQS25]. time-fractional [JCHW25, Tan25]. time-varying [LHZL25, ZZ25b]. tolerance [YZM25]. tolerant [CGH^{+}25, KNKR25]. Topological [MAMFBM^{+}25]. total [CDD^{+}25, GVW25, KSFD25]. tracking [BAT25, JYLS25, LC25, ZWQC25, ZLA25, ZZ25b]. trading [QZ25]. traffic [lCXZ^{+}25]. trajectory [KNKR25, ZLA25]. transfer [WZP^{+}25]. transform [CDM25, MH25]. transformation [SHY25]. transforms [KXL^{+}25, dOKA25]. transport [CFS25]. treatment [BDL25]. trees [BVZ25, LSWZ25, WJBL25, YT25]. triggered [DZY^{+}25, HWC25, LHZL25, LC25, QX25a, YGY^{+}25, YZ25, ZWQC25, ZXW^{+}25, ZZ25c, ZLCL25]. triggering [WZP^{+}25, YBSC25]. trust [SGH^{+}25]. tumor [BDL25]. tuning [LWS^{+}25]. Two [IAK25, LCL25, MR25, WMW^{+}25, YQYS25, BVZ25, BGQR25, LJ25, Tan25, WWS^{+}25, ZW25]. two-branch [BVZ25]. two-dimensional [LJ25, WWS^{+}25]. Two-disjoint-cycle-cover [LCL25]. two-grid [Tan25]. two-layer [BGQR25]. Two-Level [IAK25, MR25]. Two-relaxation-time [YQYS25]. type [BSG^{+}25, CDM25, Liu25, RPD25, ZZ25b]. type-2 [ZZ25b].

UAV [KNKR25]. umbral [ADZ25]. unbounded [TKM25]. uncertain [BAT25, DZY^{+}25, LWS^{+}25, XZN^{+}25]. uncertainty [CDM25]. unconstrained [YY25]. unicyclic [MP25, WJBL25]. unified [EST^{+}25]. uniform [JSZ25, PP25, Tan25, ZZ25a]. unknown [CGH^{+}25, JYLS25, QX25a, QX25b, YGY^{+}25, ZZ25d, ZLCL25]. updating [LW25b]. using [BSG^{+}25, JH25, LT25, MBF25, OGPFVV25].

vaccination [LdMAS25]. vaccine [LdMAS25]. validation [KBM25]. valuation [TLSVGR25, XZN^{+}25]. value [BGI25, HRS25]. values [CF25]. variable [QXZ25]. variants [LT25]. variational [HZF^{+}25]. Variety [KV25]. varying [LHZL25, ZZ25b]. Vector [Kov25, CFS25, ZZT25]. vehicles [lCXZ^{+}25]. verifying [TO25]. vessels [RPS^{+}25]. via [CCLL25, EST^{+}25, LW25a, WDMM25, WWS^{+}25, YGY^{+}25, YWJ^{+}25]. view [ADZ25]. viewpoint [TF25]. virtual [HAD25]. visibility [CDD^{+}25, KV25, RKS25]. volatility [Ruj25]. voltage [LC25]. Volterra [Che25, PTB25]. vs [DAG25].

walk [ZHF25]. wave [ALC^{+}25]. weak [Tan25]. weakly [KXL^{+}25, PTB25]. WeiBo [TF25]. Weighted [MWS25, GX25]. Weyl [CDM25]. wheel [ZLW25]. without [BSG^{+}25, GCV25, SHY25]. Wright [AA25]. written [XZN^{+}25].

Yeh [DS25].

Zero [ZJX25, GPL25, Ued25]. Zerodeterminant [ZJX25, Ued25]. zero-sum [GPL25]. zone [LHZL25].

References

Askari:2025:SPM

[AA25] Hassan Askari and Alireza Ansari. Stokes phenomenon for the M-Wright function of order 1n. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005496.

Aghapour:2025:DTN

[AAS25] Ahmad Aghapour, Hamid Arian, and Luis Seco. Deeptime neural networks: an efficient approach for solving high-dimensional PDEs. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005782.

Abiad:2025:DGT

[AAV25] Aida Abiad, Carlos A. Alfaro, and Ralihe R. Villagrán. Distinguishing graphs by their spectra, Smith normal forms and complements. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006593.

Araujo:2025:ITG

[ADPS25] Julio Araujo, Mitre C. Dourado, Fábio Protti, and Rudini Sampaio. The iteration time and the general position number in graph convexities. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005459.

Artioli:2025:THL

[ADZ25] M. Artioli, G. Dattoli, and U. Zainab. Theory of Hermite and Laguerre Bessel function from the umbral point of view. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005642.

An:2025:OPR

[AJK25] Jongbong An, Junkee Jeon, and Takwon Kim. Optimal portfolio and retirement decisions with costly job switching options. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006763.

Alcantara:2025:NAN

[ALC^{+}25] Adriano A. Alcântara, Juan B.

Límaco, Bruno A. Carmo, Ronald R. Guardia, and Mauro A. Rincon. Numerical analysis for nonlinear wave equations with boundary conditions: Dirichlet, acoustics and impenetrability. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004703.

Aledo:2025:MSS

[ALV25] Juan A. Aledo, Jose P.

Llano, and Jose C. Valverde. Multi-state sequential Boolean networks: an analysis of its combinatorial dynamics. Applied Mathematics and Computation, 489(??): ??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005769.

Anonymous:2025:Aa

[Ano25a] Anonymous. 1 April 2025.

Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Fa

[Ano25b] Anonymous. 1 February 2025. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ja

[Ano25c]  Anonymous.  1 January

2025. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ma

[Ano25d]  Anonymous.  1 March

2025. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ab

[Ano25e]  Anonymous.  15 April

2025. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Fb

[Ano25f] Anonymous. 15 February 2025. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jb

[Ano25g] Anonymous. 15 January 2025. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:M

[Ano25h]  Anonymous.  15 March

2025. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:EBa

[Ano25i]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005101.

Anonymous:2025:EBb

[Ano25j]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005241.

Anonymous:2025:EBc

[Ano25k]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006015.

Anonymous:2025:EBd

[Ano25l]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006313.

Anonymous:2025:EBe

[Ano25m]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006544.

Anonymous:2025:EBf

[Ano25n]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006866.

Anonymous:2025:EBg

[Ano25o]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400701X.

Anonymous:2025:EBh

[Ano25p]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007227.

Ackley:2021:LDN

[AS21] Matthew Ackley and Peter Stechlinski. Lexicographic derivatives of nonsmooth glucose-insulin kinetics under normal and artificial pancreatic responses. Applied Mathematics and Computation, 395(??):??, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300320308298.

See corrigendum [AS25].

Ackley:2025:CLD

[AS25] Matthew Ackley and Peter Stechlinski. Corrigendum to “Lexicographic derivatives of nonsmooth glucose– insulin kinetics under normal and artificial pancreatic responses” [Appl. Math. Compt. 395 (2021) 125876]. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005940.

See [AS21].

Baghi:2025:SAO

[BAT25] Hamidreza Baghi, Farzaneh Abdollahi, and Heidar Ali Talebi. Secure adaptive output-feedback tracking control for a class of uncertain nonlinear systems in the presence of sensor and actuator faults under DoS and injection attacks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400568X.

Bulai:2025:NSM

[BDL25] I. M. Bulai, M. C. De Bonis, and C. Laurita. Numerical solution of metastatic tumor growth models with treatment. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004491.

Bresar:2025:PCD

[BFH^{+}25] Bostjan Bresar, Jasmina Ferme, Premysl Holub, Marko Jakovac, and Petra Melicharová. S-packing colorings of distance graphs with distance sets of cardinality 2. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006611.

Brugnano:2025:SNP

[BGI25] Luigi Brugnano, Gianmarco Gurioli, and Felice Iavernaro. A shooting-Newton procedure for solving fractional terminal value problems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006258.

Besabe:2025:LNF

[BGQR25] Lander Besabe, Michele Girfoglio, Annalisa Quaini, and Gianluigi Rozza. Linear and nonlinear filtering for a two-layer quasi-geostrophic ocean model. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005824.

Buckwar:2025:NAC

[BM25] Evelyn Buckwar and Amira Meddah. Numerical approximations and convergence analysis of piecewise diffusion Markov processes, with application to glioma cell migration. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006945.

Bilfaqih:2025:TDR

[BQS25] Yusuf Bilfaqih, Mochamad Nur Qomarudin, and Mochammad Sahal. Time-dependent reliability computation of system with multistate components. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005435.

Bate:2025:JTM

[BSG^{+}25] Indra Bate, Kedarnath Senapati, Santhosh George, Muniyasamy M., and Chandhini G. Jarratt-type methods and their convergence analysis without using Taylor expansion. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005733.

Bantva:2025:RLT

[BVZ25] Devsi Bantva, Samir Vaidya, and Sanming Zhou. Radio labelling of two-branch trees. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005587.

Cui:2025:ESS

[CCLL25] Qian Cui, Jinde Cao, Lulu Li, and Yang Liu. Exponential stability of switched systems with state-dependent delayed impulses via Bequivalent method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004922.

Cicerone:2025:MTM

[CDD^{+}25] Serafino Cicerone, Alessia Di Fonso, Gabriele Di Stefano, Alfredo Navarra, and Francesco Piselli. Mutual and total mutual visibility in hypercube-like graphs. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006775.

Chen:2025:THP

[CDM25] Xuan Chen, Pei Dang, and Weixiong Mai. L^{p}- type Heisenberg–Pauli–Weyl uncertainty principles for fractional Fourier transform. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006970.

Cococcioni:2025:LPI

[CF25] Marco Cococcioni and Lorenzo Fiaschi. Linear programming with infinite, finite, and infinitesimal values in the right-hand side. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005058.

Chen:2025:NCG

[CFS25] Kangming Chen, Ellen Hidemi Fukuda, and Hiroyuki Sato. Nonlinear conjugate gradient method for vector optimization on Riemannian manifolds with retraction and vector transport. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004624.

Chalons:2025:SOS

[CG25] Christophe Chalons and Mathieu Girardin. Secondorder schemes based on a relaxation system for scalar and systems of conservation laws. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006349.

Cheng:2025:FTA

[CGH^{+}25] Xuanrui Cheng, Ming Gao, Wuxiang Huai, Yichun Niu, and Li Sheng. Fixedtime active fault-tolerant control for dynamical systems with intermittent faults and unknown disturbances. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005150.

Chen:2025:PDL

[Che25] Mengxin Chen. Pattern dynamics of a Lotka–Volterra model with taxis mechanism. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004788.

Chen:2025:NNI

[CjZjHbX25] Tao Chen, Yao jia Zhang, Nan jing Huang, and Yi bin Xiao. A new nonlocal impulsive fractional differential hemivariational inclusions with an application to a frictional contact problem. Applied Mathematics and Computation, 490(??): ??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006726.

Chen:2025:MEE

[CLLC25] Xiaofeng Chen, Dongyuan Lin, Hua Li, and Zhi Cheng. Minimum error entropy high-order extend Kalman filter with fiducial points. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005745.

Canovas:2025:DLH

[CMG25] J. S. Cánovas and M. Muñoz-Guillermo. On the dynamics of a linear-hyperbolic population model with Allee effect and almost sure extinction. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004661.

Chen:2025:FBD

[CWWT25] Jianjun Chen, Yuxuan Wang, Weiyi Wang, and Zhijun Tan. A FFT-based DDSIIM solver for elliptic interface problems with discontinuous coefficients on arbitrary domains and its error analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005472.

Divya:2025:MPP

[DAG25] G. Divya, S. Athithan, and Mini Ghosh. Modelling the prevalence of prostitution under the influence of poverty: a deterministic vs. stochastic approach. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004910.

deOliveira:2025:ITE

[dOKA25] André J. P. de Oliveira, Diego C. Knupp, and Luiz A. S. Abreu. Integral transforms for explicit source estimation in non-linear advection–diffusion problems. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005538.

Dou:2025:CBM

[DS25] Donna Quanjie Dou and Lisa Hui Sun. A conjecture on Boros–Moll polynomials due to Ma, Qi, Yeh and Yeh. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006477.

Deng:2025:RBR

[DWL^{+}25] Lili Deng, Rugen Wang, Ying Liao, Ronghua Xu, and Cheng Wang. The reputation-based reward mechanism promotes the evolution of fairness. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005034.

Du:2025:DTE

[DZY^{+}25] Zhaoping Du, Chi Zhang, Xiaofei Yang, Hui Ye, and Jianzhen Li. Discrete-time event-triggered h-infinity stabilization for three closedloop cyber-physical system with uncertain delay. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005885.

Elayabharath:2025:ROB

[EST^{+}25] V. T. Elayabharath, P. Sozhaeswari, N. Tatar, R. Sakthivel, and T. Satheesh. Resilient observer-based unified state and fault estimation for nonlinear parabolic PDE systems via fuzzy approach over finite-time interval. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005861.

Fome:2025:ACR

[FBK25] Anna Daniel Fome, Wolfgang Bock, and Axel Klar. Analysis of a competitive respiratory disease system with quarantine: Epidemic thresholds and crossimmunity effects. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004296.

Fermo:2025:GAM

[FLLR25] Luisa Fermo, Anna Lucia Laguardia, Concetta Laurita, and Maria Grazia Russo. A global approximation method for secondkind nonlinear integral equations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005551.

Gurubachan:2025:AFF

[GCV25] Gurubachan, V. V. M. S.

Chandramouli, and S. Verma. Analysis of α-fractal functions without boundary point conditions on the Sierpiński gasket. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005332.

Gupta:2025:SRH

[GKM25] Saloni Gupta, Arnab Kayal, and Moumita Mandal. Superconvergence results for hypersingular integral equation of first kind by Chebyshev spectral projection methods. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032400554X.

Geng:2025:ECA

[GLHS25] Yini Geng, Yikang Lu, Lijun Hong, and Lei Shi. Evolution of cooperation with asymmetric rewards.

Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005368.

Gao:2025:EEC

[GPH25] Liyan Gao, Qiuhui Pan, and Mingfeng He. The effect of ecological compensation on a coevolutionary common pool resource game. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006696.

Guo:2025:DED

[GPL25] Siyu Guo, Yingnan Pan, and Hongyi Li. Dynamic event-driven optimal consensus control for stateconstrained multiagent zerosum differential graphical games. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004405.

Gao:2025:CUE

[GSL25] Shiping Gao, Jinghui Suo, and Nan Li. Cooperation under endogenous punishment in the spatial public goods game. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006179.

Guo:2025:STR

[GVW25] Yubao Guo, Lutz Volkmann, and Yun Wang. Signed total Roman domination and domatic numbers in graphs. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005356.

Gao:2025:WSS

[GX25] Jingjing Gao and Xiangpeng Xie. A weighted switching sequence optimization algorithm for static output feedback control synthesis of nonlinear systems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006131.

Hooshyarfarzin:2025:NSE

[HAD25] Baharak Hooshyarfarzin, Mostafa Abbaszadeh, and Mehdi Dehghan. Numerical simulation and error estimation of the Davey– Stewartson equations with virtual element method. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006076.

Huang:2025:CPD

[HCCH25] Xianbin Huang, Liming Chen, Wangyong Chen, and Yao Hu. Change point detection in temporal networks based on graph snapshot similarity measures. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400626X.

Han:2025:ESM

[HRS25] Qiansheng Han, Antti Rasila, and Tommi Sottinen. Efficient simulation of mixed boundary value problems and conformal mappings. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005800.

Hou:2025:PFT

[HWC25] Xinzhao Hou, Huaiqin Wu, and Jinde Cao. Practical finite-time synchronization for Lur’e systems with performance constraint and actuator faults: a memorybased quantized dynamic event-triggered control strategy. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005691.

Han:2025:LRQ

[HYK^{+}25] Juan Han, Liqiao Yang, Kit Ian Kou, Jifei Miao, and Lizhi Liu. Low-rank quaternion matrix completion based on approximate quaternion SVD and sparse regularizer. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400691X.

Huang:2025:PBA

[HZF^{+}25] Ming Huang, Si Qi Zhang, Yong Xiu Feng, Jin Long Yuan, and Hong Han Bei. A proximal bundle approach for solving the generalized variational inequalities with inexact data. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005484.

Iqbal:2025:TLM

[IAK25] Azhar Iqbal, Shahbaz Ahmad, and Junseok Kim. Two-level method for blind image deblurring problems. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004697.

Iman:2025:DDP

[IJM25] Sara Iman and Mohammad-Reza Jahed-Motlagh. Datadriven predictor of controlaffine nonlinear dynamics: Finite discrete-time bilinear approximation of Koopman operator. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005290.

Izgi:2025:HMN

[IÖÜP25] Burhaneddin Izgi, Murat Özkaya, Nazim Kemal Üre, and Matjaz Perc. A holistic matrix norm-based alternative solution method for Markov reward games. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400585X.

Jiang:2025:FDA

[JCHW25] Yubing Jiang, Hu Chen, Chaobao Huang, and Jian Wang. A fully discrete GL-ADI scheme for 2D time-fractional reactionsubdiffusion equation. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006088.

Jarolimova:2025:DNS

[JH25] Alena Jarolímová and Jaroslav Hron. Determination of Navier’s slip parameter using data assimilation. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006180.

Jiao:2025:JSC

[JLZ25] Yujian Jiao, Tingting Li, and Zhongqiang Zhang. Jacobi spectral collocation method of space-fractional Navier–Stokes equations. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005721.

Jiang:2025:NFN

[JSZ25] Taijiang Jiang, Qiang Sun, and Chao Zhang. A note on the 1-factorization of nonuniform complete hypergraph. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004685.

Jia:2025:RLB

[JYLS25] Guolong Jia, Qing Yang, Jinxu Liu, and Hao Shen. Reinforcement learning-based linear quadratic tracking control for partially unknown Markov jump singular interconnected systems. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006908.

Kunstek:2025:MCV

[KBM25] Petar Kunstek, Martina Bukac, and Boris Muha. Mass conservation in the validation of fluid-poroelastic structure interaction solvers. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005423.

Keng:2025:CPL

[KK25a] Ying Ying Keng and Kiam Heong Kwa. Contagion probability in linear threshold model. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005514.

Kim:2025:SID

[KK25b] Taekyun Kim and Dae San Kim. Some identities on degenerate harmonic and degenerate higher-order harmonic numbers. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400506X.

Kim:2025:RCP

[KL25] Do Wan Kim and Donghwan Lee. Relaxed conditions for parameterized linear matrix inequality in the form of nested fuzzy summations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400540X.

Khodaverdian:2025:FTM

[KNKR25] Maria Khodaverdian, Majdeddin Najafi, Omid Kazemifar, and Shahabuddin Rahmanian. Fault-tolerant model predictive sliding mode control for trajectory replanning of multi-UAV formation flight. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005344.

Kovacevic:2025:VMM

[Kov25] Mladen Kovacević. Vector multispaces and multispace codes. Applied Mathematics and Computation, 486 (??):??, February 1, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005022.

Knor:2025:EAG

[KSFD25] Martin Knor, Riste Skrekovski, Slobodan Filipovski, and Darko Dimitrov. Extremizing antiregular graphs by modifying total σirregularity. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400660X.

Korze:2025:VMV

[KV25] Danilo Korze and Aleksander Vesel. Variety of mutual-visibility problems in hypercubes. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006799.

Kang:2025:FAA

[KXL^{+}25] Hongchao Kang, Chunzhi Xiang, Guidong Liu, Ao Liu, and Xinrui Hou. Fast algorithms and analysis of oscillatory and weakly singular generalized Bessel transforms. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006672.

Liu:2025:NMO

[LBLZ25] Jie Liu, Driss Boutat, Da-Yan Liu, and Xue-Feng Zhang. Nonlinear MIMO observable normal forms with output injection and output diffeomorphism. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006350.

Liu:2025:FDE

[LC25] Xin-Yu Liu and Wei-Wei Che. Fully distributed eventtriggered secondary voltage resilient tracking control for microgrids. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006453.

Li:2025:TDC

[LCL25] Hao Li, Liting Chen, and Mei Lu. Two-disjoint-cycle-cover pancyclicity of splitstar networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005460.

Li:2025:SOA

[LCLY25] Xinying Li, Chao Chang, Jianze Li, and Yuning Yang. A sparse optimization approach for simultaneous orthogonal tensor diagonalization. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006647.

Cen:2025:STF

[lCXZ^{+}25] Bing ling Cen, Yu Xue, Kun Zhang, Lisi Jia, and Hong di He. Study on traffic flows with connected vehicles and human-driven vehicles. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400643X.

Lu:2025:ISP

[LdMAS25] Yikang Lu, Alfonso de Miguel-Arribas, and Lei Shi. Influence of social peers on vaccine hesitancy under imperfect vaccination. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006751.

Lu:2025:EPR

[LDT^{+}25] Yikang Lu, Hui Dai, Huaiyu Tan, Xiaofang Duan, Lei Shi, and Junpyo Park. Enhancement of persistence in the rock–paper–scissors dynamics through higherorder interactions. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005447.

Li:2025:RLB

[LHZL25] Xin Li, Dakuo He, Qiang Zhang, and Hailong Liu. Reinforcement learning-based adaptive event-triggered control of multi-agent systems with time-varying deadzone. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005204.

Liu:2025:ECN

[Liu25] Qiong Liu. The equivalent conditions for norm of a Hilbert-type integral operator with a combination kernel and its applications. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400537X.

Li:2025:SGF

[LJ25] Jiajun Li and Ying Jiang.

Stable generalized finite element for two-dimensional and three-dimensional nonhomogeneous interface problems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005174.

Liu:2025:PEP

[LL25] Yandi Liu and Yonghui Li. A payoff equality perspective for evolutionary games: Mental accounting and cooperation promotion. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005009.

Lotstedt:2025:NIM

[LM25] Per Lötstedt and Sonja Mathias. Numerical integration of mechanical forces in center-based models for biological cell populations. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005307.

Li:2025:NAR

[LMQZ25] Rui Li, Yuede Ma, Zhongmei Qin, and Yingping Zhao. Note on the anti-Ramsey number for matching in hypercubes. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006155.

Li:2025:GBN

[LSQ25] Yinkui Li, Guiyu Shi, and Xiaoxiao Qin. Graphs with burning number three. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005617.

Liu:2025:NCE

[LSWZ25] Fenjin Liu, Ke Su, Wei Wang, and Hao Zhang. Non-cospectral equienergetic trees of diameter at most four. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005654.

Laadhari:2025:EFE

[LT25] Aymen Laadhari and Helmi Temimi. Efficient finite element strategy using enhanced high-order and second-derivative-free variants of Newton’s method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005198.

Luo:2025:LSP

[LTH^{+}25] Fangwen Luo, Qiong Tang, Yiting Huang, Yanhui Ding, and Sijia Tang. Local structure-preserving algorithms for the nonlinear Schrödinger equation with power law nonlinearity. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004478.

Liu:2025:PSS

[LW25a] Yang-Fan Liu and Huai-Ning Wu. Probabilistic stability and stabilization of humanmachine system via hidden semi-Markov modeling approach. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006143.

Lu:2025:PPD

[LW25b] Shounan Lu and Yang Wang. Past-performance-driven strategy updating promote cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006817.

Liu:2025:DOB

[LWS^{+}25] Ming-Rui Liu, Li-Bing Wu, Hong Sang, Liang-Dong Guo, and Sheng-Juan Huang. Disturbance observer-based fuzzy adaptive switched tuning control of uncertain nonlinear systems with full state constraints. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400571X.

Li:2025:NSE

[LX25] Yuying Li and Kexiang Xu. On the number of F k, 4-saturating edges. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006234.

Lv:2025:DAM

[LYD25] Yuan-Wei Lv, Guang-Hong Yang, and Georgi Marko Dimirovski. Distributed adaptive moving horizon estimation for multi-sensor networks subject to quantization effects. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005873.

Liu:2025:FDF

[LYLS25] Xuelong Liu, Guoju Ye, Wei Liu, and Fangfang Shi. Fuzzy discrete fractional granular calculus and its application to fractional cobweb models. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006374.

Li:2025:MLL

[LYM25] Dan Li, Minghui Yan, and Jixiang Meng. Minimizing the least Laplacian eigenvalue of signed complete graphs. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004636.

Li:2025:SEG

[LZ25] Shuchao Li and Zihan Zhou. On spectral extrema of graphs with given order and generalized 4-independence number. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400479X.

Molina-Abril:2025:TSF

[MAMFBM^{+}25] H. Molina-Abril, M. J.

Morón-Fernández, M. Benito-Marimón, F. Díaz del Río, and P. Real. Topological scale framework for hypergraphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004508.

Meier:2025:MFS

[MBF25] Andreas Meier, Eberhard Bänsch, and Florian Frank. Multiscale flow simulations of dilute polymeric solutions with bead-rod chains using Brownian configuration fields. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005526.

Ma:2025:NMM

[MCH25a] Y. Ma, C. S. Chen, and Y. C. Hon. A novel meshless method for solving long-term evolution problem on irregular domain. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006702.

Mei:2025:ACA

[MCH25b] Xu Mei, Jun Cheng, and Wentao Huang. Attackcompensated asynchronous output feedback control for stochastic switching systems with sojourn probability. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004855.

Ma:2025:IAO

[MCHB25] Liyang Ma, Daniel Conus, Wei-Min Huang, and Paolo Bocchini. Improved algorithm for the optimal quantization of single- and multivariate random functions. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004892.

Moon:2025:IFA

[MH25] Sunghwan Moon and Markus Haltmeier. Inversion formulas for the attenuated conical Radon transform: Plane and cylinder case. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006209.

Mu:2025:HCP

[MLTL25] Yu Mu, Wing-Cheong Lo, Yuanshun Tan, and Zijian Liu. Hybrid control for the prey in a spatial preypredator model with cooperative hunting and fear effect time lag. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006787.

Mapui:2025:LLC

[MM25] Arnab Mapui and Santwana Mukhopadhyay. Lyapunovlike conditions for prescribedtime stability of perturbed impulsive systems. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006489.

Munoz:2025:ELL

[MMNP25] Francisco J. Muñoz, Luca Meacci, Juan Carlos Nuño, and Mario Primicerio. Exploring the limits of the law of mass action in the mean field description of epidemics on Erdős–Rényi networks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004806.

Moon:2025:LED

[MP25] Sunyo Moon and Seungkook Park. Laplacian eigenvalue distribution for unicyclic graphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004831.

Mittal:2025:TLI

[MR25] Kajal Mittal and Rajendra K. Ray. Two-level implicit high-order compact scheme in exponential form for 3D quasilinear parabolic equations. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004454.

Ma:2025:WMP

[MWS25] Haifeng Ma, Wen Wang, and Predrag S. Stanimirović. Weighted Moore–Penrose inverses for dual matrices and its applications. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006064.

Orizaga:2025:SAC

[OGPFVV25] Saulo Orizaga, Gilberto González-Parra, Logan Forman, and Jesus Villegas-Villanueva. Solving Allen– Cahn equations with periodic and nonperiodic boundary conditions using mimetic finite-difference operators. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004545.

Palubeckis:2025:FLS

[Pal25] Gintaras Palubeckis. A fast local search based memetic algorithm for the parallel row ordering problem. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005010.

Palav:2025:RFO

[PP25] Mansi S. Palav and Vikas H. Pradhan. Redefined fourth order uniform hyperbolic polynomial B-splines based collocation method for solving advection-diffusion equation. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004533.

Prasina:2025:SCA

[PPN^{+}25] A. Prasina, V. Samuthira Pandi, W. Nancy, K. Thilagam, K. Veena, and A. Muniyappan. Synchronization and chimeras in asymmetrically coupled memristive Tabu learning neuron network. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324006246.

Perez:2025:LCS

[PSM25] R. Pérez, W. Sánchez, and H. J. Martínez. On the local convergence of a smoothing structured secant method for nonlinear complementarity. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005563.

Peyrovan:2025:CAC

[PTB25] P. Peyrovan, A. Tari, and H. Brunner. Convergence analysis of collocation solutions for delay Volterra integral equations with weakly singular kernels. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005836.

Qian:2025:DLP

[QHZ25] Yanxia Qian, Yunqing Huang, and Yongchao Zhang. Decoupled, linear and positivitypreserving schemes for a modified phase field crystal system incorporating longrange interactions. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005502.

Qiao:2025:DPC

[QM25] Hongwei Qiao and Jixiang Meng. Disjoint path covers of star graphs. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005599.

Qi:2025:ETS

[QX25a] Xiaojing Qi and Shengyuan Xu. Event-triggered samplingbased singularity-free fixedtime control for nonlinear systems subject to input saturation and unknown control directions. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005319.

Qiao:2025:MAS

[QX25b] Yu Qiao and Xiangtuan Xiong. A mollifier approach to the simultaneous identification of the unknown source and initial distribution in a spacefractional diffusion equation. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006362.

Qiu:2025:RKH

[QXZ25] Renjun Qiu, Ming Xu, and Pengfei Zhu. Reproducing kernel Hilbert space method for high-order linear Fredholm integro-differential equations with variable coefficients. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006222.

Qiu:2025:ITR

[QZ25] Jixiu Qiu and Yonghui Zhou.

Insider trading at a random deadline with correlation between dynamic asset and stochastic liquidity. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005812.

Raikov:2025:KPM

[RB25] I. O. Raikov and Y. M. Beltukov. The kernel polynomial method based on Jacobi polynomials. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006684.

Roy:2025:MVG

[RKS25] Dhanya Roy, Sandi Klavzar, and Aparna Lakshmanan S. Mutual-visibility and general position in double graphs and in Mycielskians. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005927.

Rublev:2025:IAA

[RPD25] G. D. Rublev, A. N. Parshikov, and S. A. Dyachkov. Improving approximation accuracy in Godunov-type smoothed particle hydrodynamics methods. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005897.

Romero:2025:RSM

[RPS^{+}25] Pau Romero, Abel Pedrós, Rafael Sebastian, Miguel Lozano, and Ignacio García-Fernández. A robust shape model for blood vessels analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005393.

Ramirez:2025:OII

[RS25] Hugo E. Ramírez and Rafael Serrano. Optimal investment with insurable background risk and nonlinear portfolio allocation frictions. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004843.

Rujivan:2025:APV

[Ruj25] Sanae Rujivan. Analytically pricing volatility options and capped/floored volatility swaps with nonlinear payoffs in discrete observation case under the Merton jumpdiffusion model driven by a nonhomogeneous Poisson process. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004909.

Su:2025:TDE

[SGH^{+}25] Xin Su, Ruqiang Guo, Shijia Hua, Linjie Liu, and Liang Zhang. Time delay effects on evolutionary dynamics of N-player trust game with diverse investment behaviors. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006386.

Sun:2025:PPB

[SHY25] Yu Sun, Cheng Hu, and Juan Yu. Pinning passivity and bipartite synchronization of fractional signed networks without gauge transformation. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005289.

Suvetha:2025:NFO

[SNP25] R. Suvetha, J. J. Nieto, and P. Prakash. Nonfragile output-feedback control for delayed memristive bidirectional associative memory neural networks against actuator failure. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400482X.

Shahab:2025:FDM

[SSH25] Muhammad Luthfi Shahab, Hadi Susanto, and Haralampos Hatzikirou. A finite difference method with symmetry properties for the highdimensional Bratu equation. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005976.

Shen:2025:RSE

[SZY25] Yi Shen, Jiemei Zhao, and Liqi Yu. Reachable set estimation of delayed secondorder memristive neural networks. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004557.

Tan:2025:SON

[Tan25] Zhijun Tan. Second-order non-uniform and fast twogrid finite element methods for non-linear timefractional mobile/immobile equations with weak regularity. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005046.

Tang:2025:AVE

[TF25] Sichen Tang and Aili Fang. Analysis of viewpoint evolution based on WeiBo data mining. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006738.

Tamimi:2025:CNS

[TGS25] Hengameh Tamimi, Mohammad Bagher Ghaemi, and Reza Saadati. Coupled nonlinear stochastic integral equations in the general form of the predatorprey model. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005848.

Tissaoui:2025:ESE

[TKM25] Yassine Tissaoui, James F. Kelly, and Simone Marras. Efficient spectral element method for the Euler equations on unbounded domains. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005411.

Trevisani:2025:MMN

[TLSVGR25] Davide Trevisani, José Germán López-Salas, Carlos Vázquez, and José Antonio García-Rodríguez. Mathematical models and numerical methods for a capital valuation adjustment (KVA) problem. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005666.

Tominec:2025:SER

[TNL25] Igor Tominec, Murtazo Nazarov, and Elisabeth Larsson. Stability estimates for radial basis function methods applied to linear scalar conservation laws. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004818.

Terao:2025:MVS

[TO25] Takeshi Terao and Katsuhisa Ozaki. Method for verifying solutions of sparse linear systems with general coefficients. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006659.

Ueda:2025:IZD

[Ued25] Masahiko Ueda. On the implementation of zerodeterminant strategies in repeated games. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324006404.

Wang:2025:GAS

[WDMM25] Xin Wang, Shihong Ding, Li Ma, and Keqi Mei. Global asymptotic stabilization via sampled-data control for a class of time-delayed systems. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006507.

Wang:2025:EHC

[WHCB25] Mei-Li Wang, Rong-Xia Hao, Jou-Ming Chang, and Sejeong Bang. 2-edge Hamiltonian connectedness:

Characterization and results in data center networks. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006581.

Wei:2025:MLI

[WJBL25] Peichao Wei, Wuding Jia, Francesco Belardo, and Muhuo Liu. Maximal Lanzhou index of trees and unicyclic graphs with prescribed diameter. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005770.

Wang:2025:TGR

[WMW^{+}25] Chen Wang, Junchi Ma, Chao Wang, Defu Lin, and Weilin Chen. Two graphs: Resolving the periodic reversibility of onedimensional finite cellular automata. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400612X.

Wang:2025:MSB

[WSX^{+}25] H.-Z. Wang, X.-C. Shangguan, D. Xiong, Y.-H. An, and L. Jin. A matrix-separation-based integral inequality for aperiodic sampled-data synchronization of delayed neural networks considering communication delay. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004934.

Wu:2025:GCC

[Wu25]  Hong-Yu Wu.  A gen-

eral class of constraint preconditioners for generalized saddle point linear systems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400609X.

Wang:2025:MES

[WWLZ25] Yuanhua Wang, Ying Wang, Haitao Li, and Wenke Zang. Matrix expressions of symmetric n-player games. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005952.

Wang:2025:NST

[WWQX25] Xin Wang, Xiaoping Wang, Haitao Qi, and Huanying Xu. Numerical simulation of time fractional Allen–Cahn equation based on Hermite neural solver. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006957.

Wei:2025:FRD

[WWS^{+}25] Jiabao Wei, Hai Wang, Kaibo Shi, Shuping He, and Chengcheng Ren. Finiteregion dissipative control for two-dimensional Roesser systems via Markov jumping mechanism. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005678.

Wang:2025:RFM

[WX25] Huaijin Wang and Chuanju Xu. On recurrence formulae of Müntz polynomials and applications. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006271.

Wang:2025:PEM

[WXD^{+}25] Junwei Wang, Weili Xiong, Feng Ding, Yihong Zhou, and Erfu Yang. Parameter estimation method for separable fractional-order Hammerstein nonlinear systems based on the on-line measurements. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005630.

Wu:2025:AEC

[WZJW25] Yifan Wu, Guangming Zhuang, Mengmeng Jiang, and Yanqian Wang. H_{∞} adaptive ETF control for degenerate jump systems under actuator faults and impulsive deception attacks with application to DCMIP device. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006192.

Wu:2025:BTC

[WZP^{+}25] Donghui Wu, Ying Zhao, Mingchuan Pang, Hong Sang, Shuanghe Yu, and Peng Wang. Bumpless transfer control of switched systems: a dynamic eventtriggering switching control approach. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006167.

Wang:2025:KIC

[WZY25] Yan Wang, Qian Zhang, and Kai Yuan. Kirchhoff indices of Cayley graphs on D_{2n} × Z_{m}. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005939.

Xu:2025:PJI

[XSB25] Wei-Ru Xu, Qian-Yu Shu, and Natália Bebiano. A pseudo-Jacobi inverse eigenvalue problem with a rankone modification. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005794.

Xin:2025:EPV

[XZN^{+}25] Yue Xin, Yi Zhang, Idin Noorani, Farshid Mehrdoust, and Jinwu Gao. Estimation of parameters and valuation of options written on multiple assets described by uncertain fractional differential equations. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005708.

Yang:2025:CFB

[YBSC25] Yu Yang, Wenshan Bi, Shuai Sui, and C. L. Philip Chen. Command-filter-based neural networks predefined time control for switched nonlinear systems with event-triggering mechanism. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006660.

Yang:2025:ETA

[YGY^{+}25] Yang Yang, Shuocong Geng, Dong Yue, Sergey Gorbachev, and Iakov Korovin. Event-triggered approximately optimized formation control of multiagent systems with unknown disturbances via simplified reinforcement learning. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006106.

Yu:2025:TRT

[YQYS25] Yuan Yu, Zuojian Qin, Haizhuan Yuan, and Shi Shu. Two-relaxation-time regularized lattice Boltzmann model for convectiondiffusion equation with spatially dependent coefficients. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005964.

Yang:2025:ERT

[YT25] Qing Yang and Yingzhi Tian. Exploring redundant trees in bipartite graphs. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004673.

Yu:2025:LBA

[YWJ^{+}25] Yaolong Yu, Zhengtian Wu, Baoping Jiang, Huaicheng Yan, and Yichen Lu. A Lagrange barrier approach for the minimum concave cost supply problem via a logarithmic descent direction algorithm. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005757.

Yamakawa:2025:CAR

[YY25] Yuya Yamakawa and Nobuo Yamashita. Convergence analysis of a regularized Newton method with generalized regularization terms for unconstrained convex optimization problems. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006805.

You:2025:GNH

[YYW^{+}25] Tao You, Linjiang Yang, Jian Wang, Peng Zhang, Jinchao Chen, and Ying Zhang. The guidance of neutral human populations maintains cooperation in the prisoner’s dilemma game. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005320.

Yang:2025:PBE

[YZ25] Meng Yang and Junyong Zhai. Predictor-based eventtriggered learning control of networked control systems with false data injection attacks and output delay. Applied Mathematics and Computation, 490(??): ??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400674X.

Yang:2025:FTA

[YZM25] Yayu Yang, Mingzu Zhang, and Jixiang Meng. Fault tolerance assessment for Hamming graphs based on rrestricted R-structure (substructure) fault pattern. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006210.

Zhang:2025:OSC

[ZCJL25] Jiannan Zhang, Ping Chen, Zhuo Jin, and Shuanming Li. Optimal strategies for collective defined contribution plans when the stock and labor markets are cointegrated. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006714.

Zhang:2025:ECA

[Zha25]  Hong Zhang.  Evolu-

tion of cooperation among fairness-seeking agents in spatial public goods game. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006441.

Zhang:2025:LEG

[ZHF25] Jiahao Zhang, Changxiang He, and Rongquan Feng. On the least eigenvalue of genuine strongly 3-walk-regular graphs. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006635.

Zhou:2025:PMC

[ZHL25] Wenjuan Zhou, Rong-Xia Hao, and Yilun Luo. Perfect matching cover indices of generalized rotation snarks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005629.

Zhang:2025:ZDS

[ZJX25] Zhipeng Zhang, Xiaotong Jiang, and Chengyi Xia. Zero-determinant strategy of finite games with implementation errors and its application into group decisionmaking. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005162.

Zhang:2025:BFB

[ZLA25] Lili Zhang, Chenglong Liu, and Liwei An. Barrierfunction based adaptive trajectory tracking control for high-order nonlinear systems with collision avoidance. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400465X.

Zhu:2025:DET

[ZLCL25] Guanchao Zhu, Min Luo, Guozeng Cui, and Ze Li. Dynamic event-triggered predefined-time adaptive attitude control for a QUAV with unknown deception attacks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004648.

Zhou:2025:ACD

[ZLHS25] Mengyu Zhou, Xingwen Liu, Qi Hu, and Feng Shu. Analysis and control of demand response in smart grids: an evolutionary game method. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005915.

Zhang:2025:CCW

[ZLW25] Guozhen Zhang, Xin Liu, and Dajin Wang. Component connectivity of wheel networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005575.

Zhang:2025:MES

[ZN25] Jinsen Zhang and Xiaobing Nie. Multiple exponential stability for short memory fractional impulsive Cohen–Grossberg neural networks with time delays. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005277.

Zhang:2025:DSL

[ZW25] Yonghong Zhang and Ligong Wang. Distance (signless) Laplacian spectra and energies of two classes of cyclic polyomino chains. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005605.

Zou:2025:IIC

[ZWF25] Lantao Zou, Yongjiang Wu, and Lihua Feng. D^{L}-integral and D^{Q}-integral n-Cayley graphs. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006398.

Zhang:2025:ETD

[ZWL25] Zidie Zhang, Daiyong Wu, and Nishan Li. Exploring threshold dynamics in a spatially heterogeneous ecosystem with memorybased diffusion and hunting cooperation on predators. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006829.

Zhai:2025:EBE

[ZWQC25] Junchang Zhai, Huanqing Wang, Yuping Qin, and Hongxia Cui. Estimationbased event-triggered fixedtime fuzzy tracking control for high-order nonlinear systems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005186.

Zhao:2025:FDS

[ZXW^{+}25] Miao Zhao, Jianxiang Xi, Le Wang, Cheng Wang, and Yuanshi Zheng. Fully distributed self-triggered secure consensus for nonlinear multiagent systems with sequential communication link scaling attacks. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006465.

Zhang:2025:BIS

[ZZ25a] Peng-Li Zhang and Xiao-Dong Zhang. Bounds for the incidence Q-spectral radius of uniform hypergraphs. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006623.

Zhao:2025:IDR

[ZZ25b] Jia-Yu Zhao and Tao Zhao.

Improving disturbance rejection and tracking performance of interval type-2 T– S fuzzy systems with timevarying delay. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006118.

Zheng:2025:ETD

[ZZ25c] Leyi Zheng and Yimin Zhou.

Event-triggered distributed consensus control of heterogeneous multi-agent system under communication and actuator faults. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005381.

Zhou:2025:ADU

[ZZ25d] Dan-Dan Zhou and Ran Zhao. Adaptive distributed unknown input observer for linear systems. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004880.

Zhang:2025:SFB

[ZZL^{+}25] Hui Zhang, Shenglong Zhou, Geoffrey Ye Li, Naihua Xiu, and Yiju Wang. A step function based recursion method for 0/1 deep neural networks. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005903.

Zhou:2025:QAS

[ZZT25] Hao Zhou, Qianqian Zhang, and Sanyi Tang. Qualitative analysis of the sliding vector field in a Filippov food chain model with integrated pest management strategy. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006490.